OpenStudy (anonymous):

What is the solution set of |x – 2| + 4 = 3

OpenStudy (anonymous):

i jus need to see how to work it out

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

|x – 2| + 4 = 3 Ix-2I=-1 NOW, x-2=-1 for x-2<0 x=1 AND x-2=1 for x-2>=0 x=3

OpenStudy (anonymous):

there is another one that i need help with too |2x + 4| < 12?

OpenStudy (anonymous):

|x – 2| + 4 = 3 actually that has no solution

OpenStudy (anonymous):

oh thats what i got

OpenStudy (anonymous):

what about this one |2x – 3| – 10 = –1? 6 or –3 6 or 3 6 or –6 No solution

OpenStudy (anonymous):

|2x – 3| – 10 = –1 I2x-3I=9 NOw, 2x-3=-9 AND 2x-3=9 x=-3 x=6

OpenStudy (anonymous):

oh i see how you got it ok thanks

OpenStudy (anonymous):

|2x + 4| < 12 -12<2x+4<12 -16<2x<8 -8<x<4

OpenStudy (anonymous):

hey what would be this one |5x – 2| greater than or equal to 8

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

i would say it is greater than

OpenStudy (anonymous):

Hint: IxI=x if x>=0 AND -x if x<0

OpenStudy (anonymous):

THAT SHOULD HELP U SOLVE THESE TYPE OF PROBLEM

OpenStudy (anonymous):

r u getting MAD

OpenStudy (anonymous):

@sauravshakya

OpenStudy (anonymous):

@85295james

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

no not mad see :)

OpenStudy (anonymous):

oh ok i dont want you guys getting mad

OpenStudy (anonymous):

medal please :)

OpenStudy (anonymous):

i gave it to somebody allready

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

then open a new question

OpenStudy (anonymous):

ok